40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 1 15:33:28 2022, Koji, Summary, BHD, Optomechanics configuration for the in-vacuum aux small optics Invac_opto_mechanics_v2.pdf
    Reply  Wed Feb 2 18:55:58 2022, Koji, Summary, BHD, Optomechanics configuration for the in-vacuum aux small optics Screen_Shot_2022-02-02_at_6.46.46_PM.png
       Reply  Wed Feb 2 21:16:23 2022, Koji, Summary, BHD, Optomechanics configuration for the in-vacuum aux small optics PXL_20220203_050156031.jpg
       Reply  Thu Apr 28 11:53:10 2022, Anchal, Summary, BHD, POP_SM4 and POP_SM5 Assembly 
          Reply  Tue May 3 14:02:09 2022, Anchal, Summary, BHD, Installed POP path in ITMX Chamber 
    Reply  Thu Apr 14 10:18:35 2022, Paco, Summary, BHD, 2 in oplev mirrors incompatible with LMR2V 
       Reply  Thu Apr 14 18:34:51 2022, Paco, Summary, BHD, ITMY Oplev reinstalled (Re: 2 in oplev mirrors incompatible with LMR2V) 
Message ID: 16780     Entry time: Thu Apr 14 18:34:51 2022     In reply to: 16778
Author: Paco 
Type: Summary 
Category: BHD 
Subject: ITMY Oplev reinstalled (Re: 2 in oplev mirrors incompatible with LMR2V) 

[Paco, Yehonathan]

We installed ITMYOL1 and ITMYOL2 on the ITMY chamber. We aligned the ITMY OpLev beam and closed the loop successfully, we then had a second round of YARM aligment, where we brought the Y peak transmission up from 0.04 counts to 0.09 counts (up by a factor of two). We still couldn't close the YARM loop but we have a better alignment.

ELOG V3.1.3-